Please solve this problem42/5° 4+j8+ j4
Answer #1

1. let 2- 4 215° z = 4+j8tj4 $17h = - Z = ki 215° - u+j8ttun 8290 2 44157- ų _(IS-909) 2 = 0.52-78) (27.0.1294 – 30. 48 20 {:

- atjb q a+jb M= (fajtor m= Vartor 0-180 – tanahan) or tanica) -a- jba -jb m= cart (-62m= Ja+ (-6) 2 0 = 180+ taille) Jo-taño

If you have any doubt please make a comment. Thank you